# IGNITE: Individualized GeNeration of Imputations in Time-series Electronic health records

### v1

@@ 1.5.2 Individualized Missingness Mask (IMM)
While multiple works investigated the importance of missingness indicators and their potential impact on prediction tasks (Sharafoddini et al.,, 2019; Che et al.,, 2018; Twala et al.,, 2008), these works did not account for individual patient-record observation frequencies. Specifically, a time-varying physiological feature that is never measured for a patient can inform the model about the patient’s state and medical needs. Moreover, the frequency of measurements can be related to the patient’s physiological state and possible deterioration. For example, multiple measurements for cardiac Troponins, sensitive biomarkers of myocardial injury, can indicate that the doctors suspect a cardiac-related diagnosis or complication  (Daubert and Jeremias,, 2010). In the same intuition, a missing measurement for the same cardiac Troponins can indicate that the patient is not suspected of developing such complications, as the clinical guidelines only recommend this measurement for diagnosing myocardial injury (Collinson,, 2022). However, this kind of feature-level missingness, can not be represented in the traditional binary missingness masks. For this purpose, we design a mask inspired by natural language processing document frequency metrics, such as term frequency-inverse document frequency (TFIDF) (Ramos et al.,, 2003) where a metric is calculated based on the relative word appearance frequency in a record. We refer to our mask as the Individualized Missingness Mask (IMM). Specifically, for each time-series feature in an individual patient record, we assign a value of 1 if the sample was observed. We formally define the mask for each time-series feature in a patient record:
@@ 1.5.4 Latent-Space Losses
Semantic loss. The semantic loss is used to ensure the learned representations are useful for performing downstream tasks. For this purpose, we concatenate the learnt representations from both VAEs and implement a classifier to predict an outcome of interest. The loss is calculated based on the cross-entropy:
@@
Contrastive loss. We also measure how close each pair of representations learnt by both VAEs for the same patient, compared to those of other patients using contrastive loss (Chopra et al.,, 2005). Given a dataset of $N$ patient records, contrastive loss’s objective is to have latent vectors derived from $\mathbf{VAE}^{\mathcal{OO}}$ and $\mathbf{VAE}^{\mathcal{IMM}}$ for the same patient to be similar to each other, yet different from those derived from different patients. Therefore the equation is based on minimizing the distance between the latent space vectors $\mathbf{z}^{\mathcal{OO}}$ and $\mathbf{z}^{\mathcal{IMM}}$ and when they correspond to the same patient and maximize the distance when the latent vectors are for other patients. For this, we use the same formulation derived by (Li et al.,, 2023):
@@ 1.5.5 Masked Imputation Task (MIT) Loss
In addition to the reconstruction loss used in VAEs, we introduce a new loss that forces the model to minimize the loss of artificially introduced missing values, which we refer to as Masked Imputation Task (MIT) loss as suggested in the work of (Du et al.,, 2023). The reconstruction loss used in calculating ELBO in the VAE component is calculated only for the observed values since the ground truth of those values cannot be accessed. The MIT loss masks a percentage of the observed values and introduces them to the model as missing values, simulating a missingess setup to force the model to generate accurate imputations. The MIT loss is calculated based on the mean squared error of the ground truth and the predicted values for the masked values.
@@ 1.5.6 Discriminator
The dual-VAE architecture is complemented with an LSTM-based classifier that outputs the probability that each sample belongs to either the real or imputed (generated) classes. Inspired by the vanilla GAN’s discriminator (Goodfellow et al.,, 2014), we refer to this classifier as the discriminator. As IGNITE generates high-quality data samples, the aim is to have the discriminator not be able to tell which ones are real and which ones are the generated imputations, implying that the imputations are as real as possible. The error produced by the discriminator is optimized along with the other the overall model’s losses.
@@
where $\gamma,\delta,\epsilon,\zeta,\theta$, and $\iota$ are scalars used to determine the ratio of each loss in the overall model’s loss, which are all finetuned on each dataset. The searched ranges of IGNITE scalar weights are shown in Supplementary Information section A. While both VAEs generate reconstructions, we’re interested in the one produced by the $VAE^{\mathcal{OO}}$, as it maps the data to the observed space and incorporates information from the IMM via the joint training with $VAE^{\mathcal{IMM}}$.
@@ 1.6 Datasets
To evaluate the proposed model we use three large-scale EHR datasets for intensive care unit patients from the United States and the Netherlands. The datasets include patients admitted to medical, surgical, and trauma ICUs, all of whom have experienced no mortality outcome during the first 48 hours. The included datasets are:
@@
    eICU, a multi-center database for ICU data, from more than 208 hospitals throughout the United States between 2014-2015, making it a good choice for validating models across multi-centers (Pollard et al.,, 2018). This database includes more than 200,000 admissions and 827 million observations. We follow the same preprocessing applied by Li et al. Li et al., (2023). The included encounters for this study are those who had at least 48 hours of ICU stay, with at least 24 hours with no mortality outcome. The final dataset included 54,423 patient encounters with 55 time-series physiological data features, as shown in Supplementary Information section B. For the treatment features, we aggregated treatments in terms of treatment concepts and kept those with at least 10% administration to patients. This resulted in three included time-series treatment features: oxygen therapy, vasopressors and antibiotics, respectively.
@@
    HiRID, a single centre critical care dataset collected from Intensive Care Medicine of the Bern University Hospital, Switzerland. The dataset contains admissions for more than 33,000 patients. Like eICU, for this study, we only included encounters for patients with a minimum of 48 hours of ICU stay and at least 24 hours with no mortality outcome. The included physiological time-series features are 50 features, shown in Supplementary Information section C. We also applied treatment concept grouping and chose those with at least 10% prevalence. The included treatments included seven features: oxygen therapy, crystalloids, vasopressors, vasodilators, insulin, painkillers and anticoagulants, respectively.
The datasets are preprocessed so that the measurements taken in each hour are aggregated, resulting in 48-time steps for each feature. If the feature was measured more than once, the arithmetic mean of hourly measurement is recorded. The value is considered missing if the feature is not measured in that respective hour. The three datasets are described in terms of missingness and included features in Table 7, Supplementary sections B and C for PhysioNet 2012, eICU and HiRID datasets, respectively. The first column refers to the overall missingness across all patient records, features and time steps, which we refer to as sample-wise missingness. The second column refers to the percentage of patients who did not have any observations of the corresponding feature, which indicates a type of missingness we refer to as feature-wise missingness.
@@ 1.7.1 Downstream Task
Due to the absence of ground-truth labels for highly missing EHRs data, we evaluate IGNITE’s imputations and compare them with those produced by the baseline methods using an extensive downstream task framework based on the original data missingness. To do so, we train an LSTM Hochreiter and Schmidhuber, (1997) classifier to predict mortality in any of the framework components explained below. The imputation models were all trained on the training set with a ratio of 80% of the full datasets and evaluated at inference by producing imputations for an unseen test set. The Supplementary Material section D shows the hyperparameter search ranges for each predictive model. The final performance is reported in terms of both area under receiving operating curve (AUROC) (Fan et al.,, 2006) and area under the precision-recall curve (AUPRC) (Sofaer et al.,, 2019). AUPRC measures the area under the precision and recall curve, making it more informative for class imbalance cases in all three datasets. In addition to the downstream task usually evaluated in imputation works, we propose an evaluation framework that tests IGNITE’s ability to impute if trained and tested on data with various types and percentages of naturally occurring missingness. The proposed downstream task framework includes:
@@
    Performance across the full population In this analysis, we evaluate the machine learning model’s performance in a mortality classification task using the full dataset, irrespective of the missingness rate. This is the common evaluation used by the related works in the literature (Fortuin et al.,, 2020; Cao et al.,, 2018; Luo et al.,, 2019, 2018).
@@
    Performance across patients with overall sample-wise missingness In this analysis, we investigate the performance of the mortality prediction classifier when trained on population subgroups where the sample-wise missingness across all features and time-steps is $\leq$ 25%, 25-75% and $\geq$ 75%, respectively. We did not report results for a subset that has less than 800 samples as they are unreliable.
@@
    Performance across patients with high feature-wise missingness While modelling the overall missingness is important, for this analysis, we focus on patients with a high percentage of features that were completely missing, i.e., features that were not observed for a particular patient. This analysis aims to evaluate the impact of various imputation methods when some features are never measured for population subgroups. Like the sample-wise missingness experiments, the investigated subgroups were also $\leq$ 25%, 25-75% and $\geq$ 75%, respectively. Similar to the sample-wise missingness experiments, the results of subsets with low-sample sizes were excluded.
@@ 1.7.2 Reconstruction Task
Despite not having the ground-truth labels for the missing values in the EHR dataset, there are observed values which we can use to evaluate the accuracy of reconstruction using the proposed model. We, therefore, randomly mask a percentage of observed values for each patient in the test set as hidden and then observe the model’s ability to recover the original values. A similar approach was used previously by (Luo et al.,, 2018; Fortuin et al.,, 2020; Nazabal et al.,, 2020). The main purpose of this validation method is to measure the models’ ability to reconstruct the original data and quantitatively compare that to other baseline methods. The results reconstructions are then compared to those of the observed values in terms of Root-Mean-Square Error (RMSE) and Mean Absolute Error (MAE). Since missingness is randomly introduced across the test set population, the RMSE is calculated for each patient and weighted by the proportion of introduced missingness in the patient record for the overall introduced missing samples. The final mean and standard deviation of RMSE and MAE across all patients are reported for each baseline. The missingness percentages used for this task were 10%, 20% and 50%, respectively.
@@ 1.7.3 Baselines
  • Transformer: An implementation of the attention is all you need paper (Vaswani et al.,, 2017), adapted for time-series imputations.
@@
To allow for the reproducibility of our work, we make our code publically available at https://github.com/Ghadeer-Ghosheh/IGNITE.
@@ Appendix D Hyper-parameters for Downstream Tasks
Appendix E Reconstruction Results for HiriD& eICU
The performance of the baseline methods in comparison to IGNITE in terms of the reconstruction task. We showcase the results for HiRID and eICU in Tables 12 and  13, respectively. In both experiments and across all percentages of masking, IGNITE consistently outperformed all the other baselines.
@@ Appendix E Reconstruction Results for HiriD& eICU
Appendix F Visualizations for Reconstructions
Below are the visualized reconstructions for two patients from the physioNet dataset. In this experiment, the trained models were used to impute randomly masked samples for each patient. The approach used here was introducing 50% missingness in the observed values for each patient record. We can see that GP-VAE seems to overfit on the training set, while LOCF imputes by zero when there’s no observed value at the start of the patient observation recording. The Visualizations for all the evaluated models are shown in 4.
(a) Imputation for a patient with sample-wise missingness.
(b) Imputation for a patient with feature-wise missingness.
Figure 4: Visualized imputations for two patients from the PhysioNet 2012 dataset. In a and b, we show examples of patients with different types of missingness. The examples shown are from the test set used for the reconstruction experiments where 50% of the observed values in the overall patient record are masked, and various imputations are compared with the ground truth. The original masked values are shown in green.

### v2

@@ 1.5.2 Individualized Missingness Mask (IMM)
Although multiple works investigated the importance of missingness indicators and their potential impact on prediction tasks (Sharafoddini et al.,, 2019; Che et al.,, 2018; Twala et al.,, 2008), these works did not account for individual patient-record observation frequencies. Specifically, a time-varying physiological feature that is never measured for a patient can inform the model about the patient’s state and medical needs. Moreover, the frequency of measurements can be related to the patient’s physiological state and possible deterioration. For example, multiple measurements for cardiac troponins, sensitive biomarkers of myocardial injury, can indicate that doctors suspect a cardiac-related diagnosis or complication  (Daubert and Jeremias,, 2010). In the same intuition, a missing measurement for the same cardiac troponins can indicate that the patient is not suspected of developing such complications, as the clinical guidelines only recommend this measurement for diagnosing myocardial injury (Collinson,, 2022). However, this kind of feature-level missingness, can not be represented in the traditional binary missingness masks. For this purpose, we design a mask inspired by natural language processing document frequency metrics, such as the term frequency-inverse document frequency (TFIDF) (Ramos et al.,, 2003) where a metric is calculated based on the relative frequency of appearance of words in a record. We refer to our mask as the Individualized Missingness Mask (IMM). Specifically, for each time-series feature in an individual patient record, we assign a value of 1 if the sample was observed. We formally define the mask for each time-series feature in a patient record:
@@ 1.5.4 Latent-Space Losses
Semantic loss. The semantic loss is used to ensure the learned representations are useful for performing downstream tasks. For this purpose, we concatenate the learned representations from both VAEs and implement a classifier to predict an outcome of interest. The loss is calculated based on the cross-entropy:
@@
Contrastive loss. We also measure how close each pair of representations learned by both VAEs for the same patient, compared to those of other patients using contrastive loss (Chopra et al.,, 2005). Given a dataset of $N$ patient records, the objective of contrastive loss is to have latent vectors derived from $\mathbf{VAE}^{\mathcal{OO}}$ and $\mathbf{VAE}^{\mathcal{IMM}}$ for the same patient to be similar to each other, yet different from those derived from different patients. Therefore the equation is based on minimizing the distance between the latent space vectors $\mathbf{z}^{\mathcal{OO}}$ and $\mathbf{z}^{\mathcal{IMM}}$ and when they correspond to the same patient and maximize the distance when the latent vectors are for other patients. For this, we use the same formulation derived by (Li et al.,, 2023):
@@ 1.5.5 Masked Imputation Task (MIT) Loss
In addition to the reconstruction loss used in VAEs, we introduce a new loss that forces the model to minimize the loss of artificially introduced missing values, which we refer to as Masked Imputation Task (MIT) loss as suggested in the work of (Du et al.,, 2023). The reconstruction loss used in calculating the ELBO in the VAE component is calculated only for the observed values, since the ground truth of those values cannot be accessed. The MIT loss masks a percentage of the observed values and introduces them to the model as missing values, simulating a missingess setup to force the model to generate accurate imputations. The MIT loss is calculated based on the mean squared error of the ground truth and the predicted values for the masked values.
@@ 1.5.6 Discriminator
The dual-VAE architecture is complemented with an LSTM-based classifier that outputs the probability that each sample belongs to either the real or imputed (generated) classes. Inspired by the vanilla GAN’s discriminator (Goodfellow et al.,, 2014), we refer to this classifier as the discriminator. As IGNITE generates high-quality data samples, the aim is to have the discriminator not be able to tell which ones are real and which ones are the generated imputations, implying that the imputations are as real as possible. The error produced by the discriminator is optimized along with the overall model’s losses.
@@
where $\gamma,\delta,\epsilon,\zeta,\theta$, and $\iota$ are scalars used to determine the ratio of each loss in the overall model’s loss, which are all finetuned on each dataset.
The optimization of IGNITE relies critically on the precise selection of hyperparameters, which was achieved by using a Bayesian hyperparameter search while minimizing the reconstruction loss. For each dataset, a customized Bayesian search was performed to accommodate its unique characteristics, such as variability in missing data patterns and feature scale. The final selection of hyperparameters was determined based on their ability to maximize imputation effectiveness and prediction accuracy, as evidenced by extensive validation experiments. This targeted approach ensured that the IGNITE framework was optimally tuned to each dataset’s specific challenges, enhancing both imputation quality and clinical utility. The searched ranges of IGNITE scalar weights are shown in the Supplementary Information Section A. While both VAEs generate reconstructions, we are interested in the one produced by the $VAE^{\mathcal{OO}}$, as it maps the data to the observed space and incorporates information from the IMM via the joint training with $VAE^{\mathcal{IMM}}$.
@@ 1.6 Datasets
To evaluate the proposed model we use three large-scale EHR datasets for intensive care unit patients from the United States and the Netherlands. The datasets include patients admitted to the ICU for medical, surgical and trauma care, all of whom have not experienced a mortality outcome during the first 48 hours. The included datasets are as follows:
@@
    eICU, a multi-center database for ICU data, from more than 208 hospitals throughout the United States between 2014-2015, making it a good choice for validating models across multi-centers (Pollard et al.,, 2018). This database includes more than 200,000 admissions and 827 million observations. We follow the same preprocessing applied by Li et al. Li et al., (2023). The included encounters for this study are those who had at least 48 hours of ICU stay, with at least 24 hours without mortality outcome. The final dataset included 54,423 patient encounters with 55 time-series physiological data features, as shown in the Supplementary Information section B. For the treatment features, we aggregated treatments in terms of treatment concepts and kept those with at least 10% administration to patients. This resulted in three included time-series treatment features: oxygen therapy, vasopressors, and antibiotics, respectively.
@@
    HiRID, a single centre critical care dataset collected from Intensive Care Medicine of the Bern University Hospital, Switzerland. The dataset contains admissions for more than 33,000 patients. Like eICU, for this study, we only included encounters for patients with a minimum of 48 hours of ICU stay and at least 24 hours without mortality outcome. The included physiological time-series features are 50 features, shown in the Supplementary Information section C. We also applied the treatment concept grouping and chose those with at least 10% prevalence. The included treatments included seven features: oxygen therapy, crystalloids, vasopressors, vasodilators, insulin, painkillers and anticoagulants, respectively.
The datasets are preprocessed so that the measurements taken in each hour are aggregated, resulting in 48-time steps for each feature. If the feature was measured more than once, the arithmetic mean of the measurement was recorded each hour. The value is considered missing if the feature is not measured in that respective hour. The three datasets are described in terms of missingness and included features in Table 7, Supplementary sections B and C for PhysioNet 2012, eICU and HiRID datasets, respectively. The first column refers to the overall missingness across all patient records, features and time steps, which we refer to as sample-wise missingness. The second column refers to the percentage of patients who did not have any observations of the corresponding feature, which indicates a type of missingness we refer to as feature-wise missingness. We formally define the two types of missingness below:
  • Feature-wise missingness: This refers to the percentage of patients who have no observations for a specific feature. It captures the extent to which a particular feature is entirely missing for a subset of patients. In this case, missingness is evaluated at the level of the feature for each patient, highlighting if some features are completely absent in certain patient records.
  • Sample-wise missingness: This refers to the overall percentage of missing values across all patient records, features, and time steps. It is calculated as the ratio of the total number of missing entries to the total number of possible entries. This metric captures how much data is missing at the level of individual data points, giving a sense of the overall missingness across the dataset.
@@ 1.7.1 Downstream Task
Due to the absence of ground-truth labels for highly missing EHRs data, we evaluate IGNITE’s imputations and compare them with those produced by the baseline methods using an extensive downstream task framework based on the original data missingness. To do so, we train an LSTM Hochreiter and Schmidhuber, (1997) classifier to predict mortality in any of the components of the framework explained below. During the development of the IGNITE model, we strictly adhered to protocols that ensure that there is no overlap or information exchange between training and test sets. The imputation model was trained exclusively on the designated training data. This data set includes comprehensive features but is entirely separate from the test dataset, which was set aside at the outset of the study. All imputation models were trained in the training set with a ratio of 80% of the full datasets and evaluated at inference by producing imputations for an unseen test set. The Supplementary Material section D shows the hyperparameter search ranges for each predictive model. The final performance is reported in terms of both the area under the receiving operating curve (AUROC) (Fan et al.,, 2006) and the area under the precision recall curve (AUPRC) (Sofaer et al.,, 2019). AUPRC measures the area under the precision and recall curve, making it more informative for class imbalance cases in all three datasets. In addition to the downstream task usually evaluated in imputation works, we propose an evaluation framework that tests IGNITE’s ability to impute if trained and tested on data with various types and percentages of naturally occurring missingness. The proposed downstream task framework includes:
@@
    Performance across the full population In this analysis, we evaluate the machine learning model’s performance in a mortality classification task using the full dataset, irrespective of the missingness rate. This is the common evaluation used by related works in the literature (Fortuin et al.,, 2020; Cao et al.,, 2018; Luo et al.,, 2019, 2018).
@@
    Performance across patients with overall sample-wise missingness In this analysis, we investigate the performance of the mortality prediction classifier when trained in population subgroups where the sample-wise missingness across all features and time-steps is $\leq$ 25%, 25-75% and $\geq$ 75%, respectively. We did not report results for a subset that has fewer than 800 samples, as they are unreliable.
@@
    Performance across patients with high feature-wise missingness While modelling overall missingness is important, for this analysis, we focus on patients with a high percentage of features that were completely missing, i.e., features that were not observed for a particular patient. This analysis aims to evaluate the impact of various imputation methods when some features are never measured for population subgroups. Like the sample-wise missingness experiments, the investigated subgroups were also $\leq$ 25%, 25-75% and $\geq$ 75%, respectively. Similarly to the sample-wise missingness experiments, the results of subsets with low-sample sizes were excluded.
@@ 1.7.2 Reconstruction Task
Despite not having the ground-truth labels for the missing values in the EHR dataset, there are observed values which we can use to evaluate the accuracy of reconstruction using the proposed model. Therefore, we randomly mask a percentage of the observed values for each patient in the test set as hidden and then observe the model’s ability to recover the original values. A similar approach was previously used by (Luo et al.,, 2018; Fortuin et al.,, 2020; Nazabal et al.,, 2020). The main purpose of this validation method is to measure the models’ ability to reconstruct the original data and quantitatively compare it to other baseline methods. In our evaluation, the quality of the imputed data is assessed by comparing the reconstructions to the observed values using Root-Mean-Square Error (RMSE) and Mean Absolute Error (MAE). To account for variability in the amount of data missing from each patient’s record, we introduce missingness randomly across the test set population. The RMSE for each patient is then calculated and weighted according to the proportion of missingness introduced in their records, ensuring that our metrics accurately reflect the severity and distribution of missing data. For this analysis, missingness was introduced at varying levels of 10%, 20%, and 50% to test the robustness of our imputation methods under different scenarios. The final results are presented as the mean and standard deviation of RMSE and MAE across all patients for each baseline method, providing a comprehensive view of performance across different degrees of data incompleteness.
@@ 1.7.3 Baselines
  • Transformer: An implementation of attention is all you need paper (Vaswani et al.,, 2017), adapted for time-series imputations.
@@
To allow the reproducibility of our work, we make our code publicly available at https://github.com/Ghadeer-Ghosheh/IGNITE.
@@ Appendix D Hyper-parameters for Downstream Tasks
Appendix E Statistical Analysis of Imputation Methods
To assess the statistical significance of the performance differences between IGNITE and other imputation methods, we applied the Wilcoxon signed-rank test. This non-parametric test is suitable for comparing paired samples and does not assume a normal distribution of the differences, making it ideal for our scenario where performance metrics may vary significantly across datasets.
We compared the AUROC and AUPRC metrics of the IGNITE model with those of six other imputation methods: LOCF, MICE, GP-VAE, Transformer, BRITS, and SAITS. The comparison was performed separately for each dataset: PhysioNet 2012, eICU, and HiRID.The following tables summarize the p-values of the Wilcoxon signed-rank test for each comparison. A p-value below 0.05 indicates a statistically significant difference favoring the IGNITE model.
Table 12: Wilcoxon Signed-Rank Test Results (P-values)
PhysioNet 2012 Dataset
Metric & Method P-value
AUROC: IGNITE vs LOCF 0.03125
AUROC: IGNITE vs MICE 0.4375
AUROC: IGNITE vs GP-VAE 0.03125
AUROC: IGNITE vs Transformer 0.625
AUROC: IGNITE vs BRITS 0.125
AUROC: IGNITE vs SAITS 0.125
AUPRC: IGNITE vs LOCF 0.03125
AUPRC: IGNITE vs MICE 0.0625
AUPRC: IGNITE vs GP-VAE 0.03125
AUPRC: IGNITE vs Transformer 0.125
AUPRC: IGNITE vs BRITS 0.125
AUPRC: IGNITE vs SAITS 0.4375
eICU Dataset
Metric & Method P-value
AUROC: IGNITE vs LOCF 0.03125
AUROC: IGNITE vs MICE 0.03125
AUROC: IGNITE vs GP-VAE 0.03125
AUROC: IGNITE vs Transformer 0.0625
AUROC: IGNITE vs BRITS 0.03125
AUROC: IGNITE vs SAITS 0.03125
AUPRC: IGNITE vs LOCF 0.03125
AUPRC: IGNITE vs MICE 0.03125
AUPRC: IGNITE vs GP-VAE 0.03125
AUPRC: IGNITE vs Transformer 0.03125
AUPRC: IGNITE vs BRITS 0.03125
AUPRC: IGNITE vs SAITS 0.03125
HiRID Dataset
Metric & Method P-value
AUROC: IGNITE vs LOCF 0.03125
AUROC: IGNITE vs MICE 0.03125
AUROC: IGNITE vs GP-VAE 0.4375
AUROC: IGNITE vs Transformer 0.0625
AUROC: IGNITE vs BRITS 0.03125
AUROC: IGNITE vs SAITS 0.0625
AUPRC: IGNITE vs LOCF 0.03125
AUPRC: IGNITE vs MICE 0.03125
AUPRC: IGNITE vs GP-VAE 0.03125
AUPRC: IGNITE vs Transformer 0.0625
AUPRC: IGNITE vs BRITS 0.4375
AUPRC: IGNITE vs SAITS 0.0625
The Wilcoxon signed-rank test results show that the IGNITE model consistently outperforms most other imputation methods, particularly in the eICU dataset, where the differences were statistically significant across all comparisons. For the PhysioNet 2012 and HiRID datasets, significant differences were observed in several comparisons, though not all. These results highlight the robustness of the IGNITE model in handling missing data and achieving superior performance across different clinical datasets.
Appendix F Reconstruction Results for HiriD& eICU
The performance of the baseline methods compared to IGNITE in terms of the reconstruction task. We present the results for HiRID and eICU in Tables 13 and  14, respectively. In both experiments and in all percentages of masking, IGNITE consistently outperformed all other baselines.
@@ Appendix F Reconstruction Results for HiriD& eICU
Appendix G Visualizations for Reconstructions
Below are the visualized reconstructions for two patients from the physioNet dataset. In this experiment, trained models were used to impute randomly masked samples for each patient. The approach used here was to introduce 50% missingness in the observed values for each patient record. We can see that GP-VAE seems to overfit the training set, while LOCF is imputed by zero when there is no observed value at the start of the patient observation recording. The visualizations for all evaluated models are shown in 5.
(1) Imputation for a patient with sample-wise missingness.
(2) Imputation for a patient with feature-wise missingness.
Figure 5: Visualized imputations for two patients from the PhysioNet 2012 dataset. In a and b, we show examples of patients with different types of missingness. The examples shown are from the test set used for reconstruction experiments where 50% of the observed values in the overall patient record are masked, and various imputations are compared with the ground truth. The original masked values are shown in green.
Appendix H Visualizations of Imputations for a dead Patient
This section provides a detailed overview of imputed values, for example, variables from dead patients, using IGNITE, BRITS and MICE models. In Figure 3, we present a focused comparison of imputed values for urine output, HCO3, and FiO2, showcasing the performance of each model. These three variables serve as key examples, with additional imputed results for the full set of patient variables included in this section. These results underscore the superior ability of IGNITE to generate clinically meaningful imputations for all patient variables.
(1) ALP
(2) Temperature
(3) Lactate
(4) WBC
(5) Na
(6) Non Invasive Diastolic Blood pressure
(7) Heart Rate
(8) Diastolic Blood Pressure
(9) pH
(10) Mg
(11) TroponinI
(12) TroponinT
(13) Bilirubin
(14) BUN
(15) Albumin
(16) AST
(17) ALT
(18) Glucose
(19) MAP
(20) K
(21) NISysABP
(22) Creatinine
(23) Cholesterol
(24) GCS
(25) NIMAP
(26) PaCo2
(27) Platlets
(28) SaO2
(29) Respiratory Rate